OpenStudy (anonymous):

You measure a football feild (100 yards long) on a detailed map and find that it is 0.5 inches long. What is the scale of the map?

OpenStudy (anonymous):

.5in=100yds .5in*1ft/12in*1yd/3ft=1/72 yd \[\frac{\frac{1}{72}yd}{100yd}=\frac{1}{720}\]

OpenStudy (anonymous):

\[\frac{\frac{1}{72}yd}{100yd}=\frac{1}{7200}\], sorry